# Towards Computation- and Communication-efficient Computational Pathology

### v1

@@ Abstract
Despite the impressive performance across a wide range of applications, current computational pathology (CPath) models face significant diagnostic efficiency challenges due to their reliance on high-magnification whole-slide image analysis. This limitation severely compromises their clinical utility, especially in time-sensitive diagnostic scenarios and situations requiring efficient data transfer. To address these issues, we present a novel computation- and communication-efficient framework called MAGnification-Aligned Global-Local Transformer (MAGA-GLTrans). Our approach significantly reduces computational time, file transfer requirements, and storage overhead by enabling effective analysis using low-magnification inputs rather than high-magnification ones. The key innovation lies in our proposed magnification alignment (MAGA) mechanism, which employs self-supervised learning to bridge the information gap between low and high magnification levels by effectively aligning their feature representations. Through extensive evaluation across various fundamental CPath tasks, MAGA-GLTrans demonstrates state-of-the-art classification performance while achieving remarkable efficiency gains: up to 10.7$\times$ reduction in computational time and over 20$\times$ reduction in file transfer and storage requirements. Furthermore, we highlight the versatility of our MAGA framework through two significant extensions: (1) its applicability as a feature extractor to enhance the efficiency of any CPath architecture, and (2) its compatibility with existing foundation models and histopathology-specific encoders, enabling them to process low-magnification inputs with minimal information loss. These advancements position MAGA-GLTrans as a particularly promising solution for time-sensitive applications, especially in the context of intraoperative frozen section diagnosis where both accuracy and efficiency are paramount.
@@ Introduction
Figure 1: Overview of the MAGA-GLTrans framework. a, Current computational pathology techniques analyze WSIs at high magnification, which poses significant diagnostic efficiency challenges. Our study introduces MAGA-GLTrans, a computationally efficient solution that processes low-magnification input images to address these limitations. b, We propose magnification alignment (MAGA) to preserve the information across different magnification levels in a self-supervised learning manner. c, MAGA aligns the embedding spaces between high- and low-magnification images, enabling fine-grained feature extraction from low-resolution inputs. d, Model Architecture of MAGA-GLTrans. It excels across various computational pathology (CPath) tasks in terms of classification performance, computational efficiency, communication efficiency, and extensibility. e, MAGA-GLTrans outperforms two mainstream WSI analysis pipelines (CLAM [16] and TransMIL [22]) across five CPath tasks. f, MAGA can be seamlessly extended to other WSI analysis architectures or histopathology-specific encoders (such as UNI [23] or CTransPath [24]), enabling high performance even with low-magnification inputs. g-h, Processing the low magnification images can greatly save the computational time (g) as well as the communication and storage overhead (h).
@@
In this article, we propose MAGA-GLTrans (Magnification-Aligned Global-Local Transformer), a computation- and communication-efficient framework designed for general-purpose WSI analysis. MAGA-GLTrans dramatically accelerates computational time from minutes to just a few seconds compared to conventional computational pathology techniques (Figure 1). Unlike traditional approaches that rely on high-resolution input images (20$\times$ or 40$\times$), MAGA-GLTrans enables the use of lower-resolution inputs (10$\times$ or 5$\times$), as illustrated in Figure 1d. By reducing the input dimension, MAGA-GLTrans not only improve the computational efficiency but also minimizes communication overhead and significantly alleviates storage burdens. While lower image resolution typically implies information loss, MAGA-GLTrans is capable of restoring information (Figure 1b-c) by aligning the low magnification features to the high magnification features, ensuring that prediction performance remains robust. We evaluate MAGA-GLTrans across six common computational pathology tasks, including cancer subtyping (TCGA-NSCLC), cancer subtyping in frozen sections (TCGA-NSCLC), prediction of three molecular events (TCGA-Glioma), and lymph node metastasis detection (CAMEYLON16). Our results demonstrate that MAGA-GLTrans achieves high-performance classification results (Figure 1e) on par with state-of-the-art models (CLAM [16] and TransMIL [22]) while reducing computational time by approximately tenfold (Figure 1g) and cutting communication and storage overhead by twentyfold (Figure 1h). Furthermore, we highlight the extensibility of MAGA (Figure 1f), showcasing its ability to enhance the efficiency of various CPath architectures and its compatibility with histopathology-specific encoders (e.g., CTransPath [24]) and foundation models (e.g., UNI [23]) to support low magnification inputs. In addition, the evaluation on a real-world clinical application, computer-assisted telepathology for intraoperative frozen section diagnosis (NSCLC subtyping on TCGA-NSCLC (frozen section)), suggests that MAGA-GLTrans is currently the optimal solution for this application due to the low computational and communication costs (Figure 6).
@@ Results
Figure 2: Classification performance, computational efficiency and storage burden (5×5$\times$5 × versus 20×20$\times$20 ×, TCGA-NSCLC, TCGA-Glioma, CAMELYON16). a-f, the classification performance of the 10-fold cross-validation experiment on three CPath tasks, including subtyping on TCGA-NSCLC (a, d), IDH mutation prediction on TCGA-Glioma (b, e) and lymph node metastasis detection on CAMELYON16 (c, f). a-c, Mean AUC of MAGA-GLTrans (5$\times$), GLTrans (5$\times$ and 20$\times$), CLAM (5$\times$ and 20$\times$) and TransMIL (5$\times$ and 20$\times$). d-f, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAGA-GLTrans (5$\times$) with two competitors at different magnification levels (5$\times$, top) and (20$\times$, bottom). g-l, Mean computational time (g, i, k) of each model and mean size of the WSI files (h, j, l) on three different datasets.
Overall performance of MAGA-GLTrans pipeline
@@
We evaluate the proposed pipeline on three publicly available datasets, including cancer subtyping on TCGA-NSCLC, molecular status predictions on TCGA-Glioma and breast cancer lymph node metastasis detection on CAMELYON16 [37]. Specifically, we predicted three important molecular events, TERT mutation, IDH mutation, and 1p19q co-deletion, on the TCGA-Glioma dataset according to the 2021 WHO Classification of Tumors of the Central Nervous System [38]. For each prediction task, we conduct quantitative comparisons with two outstanding WSI analysis approaches, CLAM [16] and TransMIL [22], at three different magnification levels (5×5$\times$5 ×, 10×10$\times$10 × and 20×20$\times$20 ×). For each magnification, we first use ImageNet pre-trained ResNet50 as the feature extractor, marked as 20×20$\times$20 ×, 10×10$\times$10 × and 5×5$\times$5 ×. For 5×5$\times$5 × and 10×10$\times$10 × magnifications, we ‘upgrade’ the extracted features by MAGA as 5×5$\times$5 × (w/ MAGA) and 10×10$\times$10 × (w/ MAGA). The patches at different magnification levels share exactly the same image contents but different image resolutions. 10-fold cross-validation is conducted on all the prediction tasks to demonstrate the overall performance. On the TCGA-Glioma and TCGA-NSCLC datasets, all the patients are randomly divided into a training set (80%), a validation set (10%), and a test set (10%). Multiple slides that come from one patient are all sampled into the same set to avoid the data leakage problem. On the CAMELYON16 dataset, since the challenge split the entire dataset into a training set and a test set. The test set remains unchanged, and the original training data is further divided into a training set (80%) and a validation set (20%). Besides the classification performance, we also record the runtime of each dataset to demonstrate the computational efficiency. Each dataset is saved at different magnifications to evaluate the storage burden. Additional details regarding the performance of 10×10$\times$10 × versus 20×20$\times$20 × are provided in Extended Figure 1. The performance of TERT mutation and 1p19q codeletion predictions on TCGA-Glioma are provided in Extended Figure 2 and Extended Figure 3, respectively. The performance of each fold, including AUC, accuracy and F1-score, in different magnification levels is provided in Supplementary Table 1 (5×5$\times$5 ×), Supplementary Table 2 (10×10$\times$10 ×) and Supplementary Table 3 (20×20$\times$20 ×).
@@
Overall, the classification performance on three datasets is demonstrated in Figure 2a-c. From the area under the receiver operating characteristic (AUC) results on the TCGA-NSCLC task, we observe that TransMIL and CLAM given 20×20$\times$20 × magnification inputs, defined as TransMIL (20×20$\times$20 ×) and CLAM (20×20$\times$20 ×), achieved a favorable AUC value of 0.9217 (95% confidence interval (CI) 0.9108-0.9325) and 0.9314 (95% CI 0.9125-0.9527), respectively. The computational time per slide is 33.35 seconds (s) for CLAM and 33.23s for TransMIL, while the storage burden per slide is around 0.5 gigabytes (GB). Given 5×5$\times$5 × inputs, the storage burden and the computational time greatly reduce to around 0.03 GB and 4.3s per slide. Unfortunately, due to the information loss in the low magnification images, the classification performance also drastically decreases to 0.6781 (95% CI 0.5981-0.7580, -23.90%) for TransMIL and 0.7746 (95% CI 0.7455-0.8037, -18.68%) for CLAM. Thanks to the self-supervised magnification alignment strategy proposed in our study, MAGA-GLTrans (5×5$\times$5 ×) is able to bridge the information gap between two magnifications, leading to an exceptional improvement with respect to AUC compared with TransMIL (5×5$\times$5 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.6781 (95% CI 0.5981-0.7580), $P<0.001$) and CLAM (5×5$\times$5 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.7746 (95% CI 0.7455-0.8037), $P<0.001$). The computational time per slide of MAGA-GLTrans (5×5$\times$5 ×) is further reduced compared with TransMIL (5×5$\times$5 ×) and CLAM (5×5$\times$5 ×) (3.15s versus 4.33s versus 4.34s). Additionally, MAGA-GLTrans (5×5$\times$5 ×) even marginally outperform TransMIL (20×20$\times$20 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.9217 (95% CI 0.9108-0.9325), $P<0.001$) and CLAM (20×20$\times$20 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.9314 (95% CI 0.9125-0.9527), $P<0.001$). To evaluate the robustness of the models, we randomly resample the patients in each fold for 500 times. The box-plot of each fold in the 10-fold cross-validation demonstrates that MAGA-GLTrans (5×5$\times$5 ×) consistently outperforms TransMIL and CLAM at 5×5$\times$5 × (Figure 1d (top)) and 20×20$\times$20 × (Figure 1d (bottom)).
@@
The same observations are found in TCGA-Glioma (IDH) and CAMELYON16. In the TCGA-Glioma (IDH) task, MAGA-GLTrans (5×5$\times$5 ×) achieves slightly higher mean AUC compared with TransMIL (20×20$\times$20 ×) (0.9312 (95% CI 0.9098-0.9525) versus 0.9246 (95% CI 0.9113-0.9380), $P<0.001$) and CLAM (20×20$\times$20 ×) (0.9312 (95% CI 0.9098-0.9525) versus 0.9155 (95% CI 0.9016-0.9295), $P<0.001$), while drastically reduces computational time (3.39s versus 24.73s versus 25.08s) and storage burden from 0.44GB to 0.03GB. In the CAMELYON16 task, MAGA-GLTrans (5×5$\times$5 ×) achieves comparable results with TransMIL (20×20$\times$20 ×) (0.8938 (95% CI 0.8820-0.9056) versus 0.8980 (95% CI 0.8822-0.9137), $P<0.001$) and CLAM (20×20$\times$20 ×) (0.8938 (95% CI 0.8820-0.9056) versus 0.8876 (95% CI 0.8676-0.9075), $P<0.001$), while with much less computational time (3.8s versus 26.19s versus 26.88s), lower storage burden (0.05GB (5×5$\times$5 ×) versus 1.32GB (20×20$\times$20 ×)).
@@
Furthermore, we conduct the same experiment at the 10×10$\times$10 × magnification, shown in Extended Figure 1. For example in the CAMELYON16 task, MAGA-GLTrans (10×10$\times$10 ×) also achieves the highest mean AUC values compared with TransMIL and CLAM in two magnification levels (5×5$\times$5 × and 20×20$\times$20 ×). It is not surprising that MAGA-GLTrans (10×10$\times$10 ×) slightly outperforms MAGA-GLTrans (5×5$\times$5 ×) (0.9108 (95% CI 0.9048-0.9168) versus 0.8938 (95% CI 0.8820-0.9056), $P<0.001$), since 10×10$\times$10 × images lose less information than 5×5$\times$5 × ones. But the computational time (8.97s versus 3.8s) and storage burden (0.17GB versus 0.05GB) at 10×10$\times$10 × magnification are higher than the ones at 5×5$\times$5 × magnification.
@@
MAGA can be extended to other WSI analysis architectures
@@
Figure 3: Extensibility of MAGA to other WSI analysis architectures. MAGA can be applied to any general-purpose WSI analysis architecture to improve computational efficiency. a-j, The plots demonstrate the performance (left) and the computational time (right) of TransMIL (a-e) and CLAM (f-j) at different magnification levels on five CPath tasks. The bars on the left represent the mean AUCs of the 10-fold cross-validation at 5×5$\times$5 ×, 10×10$\times$10 × and 20×20$\times$20 × with MAGA (dash border) and without MAGA. The bars on the right show the computational time at different magnification levels. Error bars represent 95% confidence intervals. The performance increment introduced by MAGA is shown on top of the performance bars.
@@
Amid the era of deep learning, the proliferation of AI technologies has greatly shortened the life cycles for algorithmic innovations and upgrades. As a result, the extensibility of models has become increasingly critical. In this experiment, we explore the extensibility of our proposed MAGA to enhance computational efficiency and communication efficiency for the other WSI analysis architectures, including TransMIL and CLAM. We perform this experiment on five CPath tasks with 10-fold cross-validation, including NSCLC subtyping (TCGA-NSCLC), three molecular status predictions (TCGA-Glioma) and breast cancer lymph node metastasis detection (CAMELYON16). All the experiments are conducted at different input magnification levels (5×5$\times$5 ×, 10×10$\times$10 × and 20×20$\times$20 ×). ImageNet pre-trained ResNet50 is used as the feature extractor for 20×20$\times$20 × magnification. When training TransMIL and CLAM at low magnification (5×5$\times$5 × and 10×10$\times$10 ×), the original pipelines remain unchanged except for replacing the feature extractor with our proposed MAGA. For each magnification level, a unique MAGA model is trained independently.
@@
We observe that the classification performance of both models given low magnification inputs significantly drops in all five CPath tasks when features are extracted using the original pipeline without MAGA. For example, the mean AUC values of TransMIL in the NSCLC subtyping task drop from 0.9314 (20×20$\times$20 ×) to 0.8274 (10×10$\times$10 ×) and 0.7746 (5×5$\times$5 ×) shown in Figure 3 (left). When equipped with MAGA, the results with low magnification inputs (5×5$\times$5 ×) and (10×10$\times$10 ×) returns to a reasonable level ($0.7746\rightarrow 0.9307$ at 5×5$\times$5 ×, $P<0.001$, $0.8274\rightarrow 0.9362$ at 10×10$\times$10 ×, $P<0.001$), which is comparable with the model trained with high magnification inputs (20×20$\times$20 ×). Overall, TransMIL and CLAM equipped with MAGA consistently outperform the models without MAGA, with the exceptional improvement of mean AUC values of (TCGA-NSCLC: $+24.56\%$ [5×5$\times$5 ×], $+8.19\%$ [10×10$\times$10 ×]), (TCGA-Glioma (IDH): $+6.81\%$ [5×5$\times$5 ×], $+5.45\%$ [10×10$\times$10 ×]), (TCGA-Glioma (TERT): $+18.78\%$ [5×5$\times$5 ×], $+5.62\%$ [10×10$\times$10 ×]), (TCGA-Glioma (1p19q): $+15.69\%$ [5×5$\times$5 ×], $+4.73\%$ [10×10$\times$10 ×]) and (CAMELYON16: $+31.43\%$ [5×5$\times$5 ×], $+10.38\%$ [10×10$\times$10 ×]) for TransMIL. MAGA also empowers CLAM with the improvement of mean AUC values of (TCGA-NSCLC: $+15.61\%$ [5×5$\times$5 ×], $+10.88\%$ [10×10$\times$10 ×]), (TCGA-Glioma (IDH): $+28.12\%$ [5×5$\times$5 ×], $+14.17\%$ [10×10$\times$10 ×]), (TCGA-Glioma (TERT): $+8.14\%$ [5×5$\times$5 ×], $+2.71\%$ [10×10$\times$10 ×]), (TCGA-Glioma (1p19q): $+11.49\%$ [5×5$\times$5 ×], $+10.17\%$ [10×10$\times$10 ×]) and (CAMELYON16: $+16.04\%$ [5×5$\times$5 ×], $+12.86\%$ [10×10$\times$10 ×]). This suggests that MAGA bridges the information gap and successfully recovers the loss information. Furthermore, we also observe a larger vibration of the performance at 5×5$\times$5 × magnification (without MAGA) than that at 10×10$\times$10 × magnification (without MAGA) due to the different degrees of information loss. Therefore, the increment of the classification performance at 5×5$\times$5 × magnification that MAGA introduces is much larger than the increment at 10×10$\times$10 × magnification. Another observation is that MAGA improves data efficiency when with small-scale data. For example, in CAMELYON16 with only 270 WSIs in the training set, the performance increment for both CLAM and TransMIL is generally larger than the other two datasets with around one thousand WSIs for model training. By maintaining a robust classification performance at low magnification levels, MAGA brings all the benefits to any WSI analysis pipeline, including computational efficiency, communication efficiency, and a storage-friendly nature. Specifically, it greatly reduces the computational time from around half a minute to only a few seconds, for example from 33.13s to 3.78s for TransMIL in TCGA-NSCLC shown in Figure 3a (right). Making current and future WSI analysis pipelines feasible and friendly to the clinical scenarios with a demand for fast diagnosis or efficient data transfer, like intraoperative diagnosis or telepathology.

### v2

@@ Abstract
Despite the impressive performance across a wide range of applications, current computational pathology (CPath) models face significant diagnostic efficiency challenges due to their reliance on high-magnification whole-slide image analysis. This limitation severely compromises their clinical utility, especially in time-sensitive diagnostic scenarios and situations requiring efficient data transfer. To address these issues, we present a novel computation- and communication-efficient framework called MAGnification-Aligned Global-Local Transformer (MAG-GLTrans). Our approach significantly reduces computational time, file transfer requirements, and storage overhead by enabling effective analysis using low-magnification inputs rather than high-magnification ones. The key innovation lies in our proposed magnification alignment (MAG) mechanism, which employs self-supervised learning to bridge the information gap between low and high magnification levels by effectively aligning their feature representations. Through extensive evaluation across various fundamental CPath tasks, MAG-GLTrans demonstrates state-of-the-art classification performance while achieving remarkable efficiency gains: up to 10.7$\times$ reduction in computational time and over 20$\times$ reduction in file transfer and storage requirements. Furthermore, we highlight the versatility of our MAG framework through two significant extensions: (1) its applicability as a feature extractor to enhance the efficiency of any CPath architecture, and (2) its compatibility with existing foundation models and histopathology-specific encoders, enabling them to process low-magnification inputs with minimal information loss. These advancements position MAG-GLTrans as a particularly promising solution for time-sensitive applications, especially in the context of intraoperative frozen section diagnosis where both accuracy and efficiency are paramount.
@@ Introduction
Figure 1: Overview of the MAG-GLTrans framework. a, Current computational pathology techniques analyze WSIs at high magnification, which poses significant diagnostic efficiency challenges. Our study introduces MAG-GLTrans, a computationally efficient solution that processes low-magnification input images to address these limitations. b, We propose magnification alignment (MAG) to preserve the information across different magnification levels in a self-supervised learning manner. c, MAG aligns the embedding spaces between high- and low-magnification images, enabling fine-grained feature extraction from low-resolution inputs. d, Model Architecture of MAG-GLTrans. It excels across various computational pathology (CPath) tasks in terms of classification performance, computational efficiency, communication efficiency, and extensibility. e, MAG-GLTrans outperforms two mainstream WSI analysis pipelines (CLAM [16] and TransMIL [22]) across five CPath tasks. f, MAG can be seamlessly extended to other WSI analysis architectures or histopathology-specific encoders (such as UNI [23] or CTransPath [24]), enabling high performance even with low-magnification inputs. g-h, Processing the low magnification images can greatly save the computational time (g) as well as the communication and storage overhead (h).
@@
In this article, we propose MAG-GLTrans (Magnification-Aligned Global-Local Transformer), a computation- and communication-efficient framework designed for general-purpose WSI analysis. MAG-GLTrans dramatically accelerates computational time from minutes to just a few seconds compared to conventional computational pathology techniques (Figure 1). Unlike traditional approaches that rely on high-resolution input images (20$\times$ or 40$\times$), MAG-GLTrans enables the use of lower-resolution inputs (10$\times$ or 5$\times$), as illustrated in Figure 1d. By reducing the input dimension, MAG-GLTrans not only improve the computational efficiency but also minimizes communication overhead and significantly alleviates storage burdens. While lower image resolution typically implies information loss, MAG-GLTrans is capable of restoring information (Figure 1b-c) by aligning the low magnification features to the high magnification features, ensuring that prediction performance remains robust. We evaluate MAG-GLTrans across six common computational pathology tasks, including cancer subtyping (TCGA-NSCLC), cancer subtyping in frozen sections (TCGA-NSCLC), prediction of three molecular events (TCGA-Glioma), and lymph node metastasis detection (CAMEYLON16). Our results demonstrate that MAG-GLTrans achieves high-performance classification results (Figure 1e) on par with state-of-the-art models (CLAM [16] and TransMIL [22]) while reducing computational time by approximately tenfold (Figure 1g) and cutting communication and storage overhead by twentyfold (Figure 1h). Furthermore, we highlight the extensibility of MAG (Figure 1f), showcasing its ability to enhance the efficiency of various CPath architectures and its compatibility with histopathology-specific encoders (e.g., CTransPath [24]) and foundation models (e.g., UNI [23]) to support low magnification inputs. In addition, the evaluation on a real-world clinical application, computer-assisted telepathology for intraoperative frozen section diagnosis (NSCLC subtyping on TCGA-NSCLC (frozen section)), suggests that MAG-GLTrans is currently the optimal solution for this application due to the low computational and communication costs (Figure 6).
@@ Results
Figure 2: Classification performance, computational efficiency and storage burden (5×5$\times$5 × versus 20×20$\times$20 ×, TCGA-NSCLC, TCGA-Glioma, CAMELYON16). a-f, the classification performance of the 10-fold cross-validation experiment on three CPath tasks, including subtyping on TCGA-NSCLC (a, d), IDH mutation prediction on TCGA-Glioma (b, e) and lymph node metastasis detection on CAMELYON16 (c, f). a-c, Mean AUC of MAG-GLTrans (5$\times$), GLTrans (5$\times$ and 20$\times$), CLAM (5$\times$ and 20$\times$) and TransMIL (5$\times$ and 20$\times$). d-f, Box plots demonstrate the results of each fold in the 10-fold cross-validation by resampling 500 times. We compare MAG-GLTrans (5$\times$) with two competitors at different magnification levels (5$\times$, top) and (20$\times$, bottom). g-l, Mean computational time (g, i, k) of each model and mean size of the WSI files (h, j, l) on three different datasets.
Overall performance of MAG-GLTrans pipeline
@@
We evaluate the proposed pipeline on three publicly available datasets, including cancer subtyping on TCGA-NSCLC, molecular status predictions on TCGA-Glioma and breast cancer lymph node metastasis detection on CAMELYON16 [37]. Specifically, we predicted three important molecular events, TERT mutation, IDH mutation, and 1p19q co-deletion, on the TCGA-Glioma dataset according to the 2021 WHO Classification of Tumors of the Central Nervous System [38]. For each prediction task, we conduct quantitative comparisons with two outstanding WSI analysis approaches, CLAM [16] and TransMIL [22], at three different magnification levels (5×5$\times$5 ×, 10×10$\times$10 × and 20×20$\times$20 ×). For each magnification, we first use ImageNet pre-trained ResNet50 as the feature extractor, marked as 20×20$\times$20 ×, 10×10$\times$10 × and 5×5$\times$5 ×. For 5×5$\times$5 × and 10×10$\times$10 × magnifications, we ‘upgrade’ the extracted features by MAG as 5×5$\times$5 × (w/ MAG) and 10×10$\times$10 × (w/ MAG). The patches at different magnification levels share exactly the same image contents but different image resolutions. 10-fold cross-validation is conducted on all the prediction tasks to demonstrate the overall performance. On the TCGA-Glioma and TCGA-NSCLC datasets, all the patients are randomly divided into a training set (80%), a validation set (10%), and a test set (10%). Multiple slides that come from one patient are all sampled into the same set to avoid the data leakage problem. On the CAMELYON16 dataset, since the challenge split the entire dataset into a training set and a test set. The test set remains unchanged, and the original training data is further divided into a training set (80%) and a validation set (20%). Besides the classification performance, we also record the runtime of each dataset to demonstrate the computational efficiency. Each dataset is saved at different magnifications to evaluate the storage burden. Additional details regarding the performance of 10×10$\times$10 × versus 20×20$\times$20 × are provided in Extended Figure 1. The performance of TERT mutation and 1p19q codeletion predictions on TCGA-Glioma are provided in Extended Figure 2 and Extended Figure 3, respectively. The performance of each fold, including AUC, accuracy and F1-score, in different magnification levels is provided in Supplementary materials.
@@
Overall, the classification performance on three datasets is demonstrated in Figure 2a-c. From the area under the receiver operating characteristic (AUC) results on the TCGA-NSCLC task, we observe that TransMIL and CLAM given 20×20$\times$20 × magnification inputs, defined as TransMIL (20×20$\times$20 ×) and CLAM (20×20$\times$20 ×), achieved a favorable AUC value of 0.9217 (95% confidence interval (CI) 0.9108-0.9325) and 0.9314 (95% CI 0.9125-0.9527), respectively. The computational time per slide is 33.35 seconds (s) for CLAM and 33.23s for TransMIL, while the storage burden per slide is around 0.5 gigabytes (GB). Given 5×5$\times$5 × inputs, the storage burden and the computational time greatly reduce to around 0.03 GB and 4.3s per slide. Unfortunately, due to the information loss in the low magnification images, the classification performance also drastically decreases to 0.6781 (95% CI 0.5981-0.7580, -23.90%) for TransMIL and 0.7746 (95% CI 0.7455-0.8037, -18.68%) for CLAM. Thanks to the self-supervised magnification alignment strategy proposed in our study, MAG-GLTrans (5×5$\times$5 ×) is able to bridge the information gap between two magnifications, leading to an exceptional improvement with respect to AUC compared with TransMIL (5×5$\times$5 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.6781 (95% CI 0.5981-0.7580), $P<0.001$) and CLAM (5×5$\times$5 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.7746 (95% CI 0.7455-0.8037), $P<0.001$). The computational time per slide of MAG-GLTrans (5×5$\times$5 ×) is further reduced compared with TransMIL (5×5$\times$5 ×) and CLAM (5×5$\times$5 ×) (3.15s versus 4.33s versus 4.34s). Additionally, MAG-GLTrans (5×5$\times$5 ×) even marginally outperform TransMIL (20×20$\times$20 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.9217 (95% CI 0.9108-0.9325), $P<0.001$) and CLAM (20×20$\times$20 ×) (0.9511 (95% CI 0.9430-0.9592) versus 0.9314 (95% CI 0.9125-0.9527), $P<0.001$). To evaluate the robustness of the models, we randomly resample the patients in each fold for 500 times. The box-plot of each fold in the 10-fold cross-validation demonstrates that MAG-GLTrans (5×5$\times$5 ×) consistently outperforms TransMIL and CLAM at 5×5$\times$5 × (Figure 1d (top)) and 20×20$\times$20 × (Figure 1d (bottom)).
@@
The same observations are found in TCGA-Glioma (IDH) and CAMELYON16. In the TCGA-Glioma (IDH) task, MAG-GLTrans (5×5$\times$5 ×) achieves slightly higher mean AUC compared with TransMIL (20×20$\times$20 ×) (0.9312 (95% CI 0.9098-0.9525) versus 0.9246 (95% CI 0.9113-0.9380), $P<0.001$) and CLAM (20×20$\times$20 ×) (0.9312 (95% CI 0.9098-0.9525) versus 0.9155 (95% CI 0.9016-0.9295), $P<0.001$), while drastically reduces computational time (3.39s versus 24.73s versus 25.08s) and storage burden from 0.44GB to 0.03GB. In the CAMELYON16 task, MAG-GLTrans (5×5$\times$5 ×) achieves comparable results with TransMIL (20×20$\times$20 ×) (0.8938 (95% CI 0.8820-0.9056) versus 0.8980 (95% CI 0.8822-0.9137), $P<0.001$) and CLAM (20×20$\times$20 ×) (0.8938 (95% CI 0.8820-0.9056) versus 0.8876 (95% CI 0.8676-0.9075), $P<0.001$), while with much less computational time (3.8s versus 26.19s versus 26.88s), lower storage burden (0.05GB (5×5$\times$5 ×) versus 1.32GB (20×20$\times$20 ×)).
@@
Furthermore, we conduct the same experiment at the 10×10$\times$10 × magnification, shown in Extended Figure 1. For example in the CAMELYON16 task, MAG-GLTrans (10×10$\times$10 ×) also achieves the highest mean AUC values compared with TransMIL and CLAM in two magnification levels (5×5$\times$5 × and 20×20$\times$20 ×). It is not surprising that MAG-GLTrans (10×10$\times$10 ×) slightly outperforms MAG-GLTrans (5×5$\times$5 ×) (0.9108 (95% CI 0.9048-0.9168) versus 0.8938 (95% CI 0.8820-0.9056), $P<0.001$), since 10×10$\times$10 × images lose less information than 5×5$\times$5 × ones. But the computational time (8.97s versus 3.8s) and storage burden (0.17GB versus 0.05GB) at 10×10$\times$10 × magnification are higher than the ones at 5×5$\times$5 × magnification.
@@
MAG can be extended to other WSI analysis architectures
@@
Figure 3: Extensibility of MAG to other WSI analysis architectures. MAG can be applied to any general-purpose WSI analysis architecture to improve computational efficiency. a-j, The plots demonstrate the performance (left) and the computational time (right) of TransMIL (a-e) and CLAM (f-j) at different magnification levels on five CPath tasks. The bars on the left represent the mean AUCs of the 10-fold cross-validation at 5×5$\times$5 ×, 10×10$\times$10 × and 20×20$\times$20 × with MAG (dash border) and without MAG. The bars on the right show the computational time at different magnification levels. Error bars represent 95% confidence intervals. The performance increment introduced by MAG is shown on top of the performance bars.
@@
Amid the era of deep learning, the proliferation of AI technologies has greatly shortened the life cycles for algorithmic innovations and upgrades. As a result, the extensibility of models has become increasingly critical. In this experiment, we explore the extensibility of our proposed MAG to enhance computational efficiency and communication efficiency for the other WSI analysis architectures, including TransMIL and CLAM. We perform this experiment on five CPath tasks with 10-fold cross-validation, including NSCLC subtyping (TCGA-NSCLC), three molecular status predictions (TCGA-Glioma) and breast cancer lymph node metastasis detection (CAMELYON16). All the experiments are conducted at different input magnification levels (5×5$\times$5 ×, 10×10$\times$10 × and 20×20$\times$20 ×). ImageNet pre-trained ResNet50 is used as the feature extractor for 20×20$\times$20 × magnification. When training TransMIL and CLAM at low magnification (5×5$\times$5 × and 10×10$\times$10 ×), the original pipelines remain unchanged except for replacing the feature extractor with our proposed MAG. For each magnification level, a unique MAG model is trained independently.
@@
We observe that the classification performance of both models given low magnification inputs significantly drops in all five CPath tasks when features are extracted using the original pipeline without MAG. For example, the mean AUC values of TransMIL in the NSCLC subtyping task drop from 0.9314 (20×20$\times$20 ×) to 0.8274 (10×10$\times$10 ×) and 0.7746 (5×5$\times$5 ×) shown in Figure 3 (left). When equipped with MAG, the results with low magnification inputs (5×5$\times$5 ×) and (10×10$\times$10 ×) returns to a reasonable level ($0.7746\rightarrow 0.9307$ at 5×5$\times$5 ×, $P<0.001$, $0.8274\rightarrow 0.9362$ at 10×10$\times$10 ×, $P<0.001$), which is comparable with the model trained with high magnification inputs (20×20$\times$20 ×). Overall, TransMIL and CLAM equipped with MAG consistently outperform the models without MAG, with the exceptional improvement of mean AUC values of (TCGA-NSCLC: $+24.56\%$ [5×5$\times$5 ×], $+8.19\%$ [10×10$\times$10 ×]), (TCGA-Glioma (IDH): $+6.81\%$ [5×5$\times$5 ×], $+5.45\%$ [10×10$\times$10 ×]), (TCGA-Glioma (TERT): $+18.78\%$ [5×5$\times$5 ×], $+5.62\%$ [10×10$\times$10 ×]), (TCGA-Glioma (1p19q): $+15.69\%$ [5×5$\times$5 ×], $+4.73\%$ [10×10$\times$10 ×]) and (CAMELYON16: $+31.43\%$ [5×5$\times$5 ×], $+10.38\%$ [10×10$\times$10 ×]) for TransMIL. MAG also empowers CLAM with the improvement of mean AUC values of (TCGA-NSCLC: $+15.61\%$ [5×5$\times$5 ×], $+10.88\%$ [10×10$\times$10 ×]), (TCGA-Glioma (IDH): $+28.12\%$ [5×5$\times$5 ×], $+14.17\%$ [10×10$\times$10 ×]), (TCGA-Glioma (TERT): $+8.14\%$ [5×5$\times$5 ×], $+2.71\%$ [10×10$\times$10 ×]), (TCGA-Glioma (1p19q): $+11.49\%$ [5×5$\times$5 ×], $+10.17\%$ [10×10$\times$10 ×]) and (CAMELYON16: $+16.04\%$ [5×5$\times$5 ×], $+12.86\%$ [10×10$\times$10 ×]). This suggests that MAG bridges the information gap and successfully recovers the loss information. Furthermore, we also observe a larger vibration of the performance at 5×5$\times$5 × magnification (without MAG) than that at 10×10$\times$10 × magnification (without MAG) due to the different degrees of information loss. Therefore, the increment of the classification performance at 5×5$\times$5 × magnification that MAG introduces is much larger than the increment at 10×10$\times$10 × magnification. Another observation is that MAG improves data efficiency when with small-scale data. For example, in CAMELYON16 with only 270 WSIs in the training set, the performance increment for both CLAM and TransMIL is generally larger than the other two datasets with around one thousand WSIs for model training. By maintaining a robust classification performance at low magnification levels, MAG brings all the benefits to any WSI analysis pipeline, including computational efficiency, communication efficiency, and a storage-friendly nature. Specifically, it greatly reduces the computational time from around half a minute to only a few seconds, for example from 33.13s to 3.78s for TransMIL in TCGA-NSCLC shown in Figure 3a (right). Making current and future WSI analysis pipelines feasible and friendly to the clinical scenarios with a demand for fast diagnosis or efficient data transfer, like intraoperative diagnosis or telepathology.